OpenStudy (anonymous):

If the center of a circle is (0, 0) and the radius is the square root of 7 , the equation of the circle is: x^2+y^2=7. Is the answer true?

OpenStudy (anonymous):

Very true.. :)

OpenStudy (anonymous):

Great, thank you. Also, for this equation of a circle (x-2)^2+(y+3)^2=6 the radius of the circle is square root of 6 or is it 36?

OpenStudy (anonymous):

sqrt{6}

OpenStudy (anonymous):

Thank you very much @waterineyes :)

OpenStudy (anonymous):

Circle equation is: \[(x-h)^2 + (y-k)^2 = r^2\] Where r is the radius.. So, comparing : \[r^2 = 6 \implies \color{blue}{r = \sqrt{6}}\]